Problem 4-A point charge q 9.0 nC is placed inside a spherical shell made of metal, of inner radius a 3.0 m and outer radius b -5.0 m, that is charged with qp-6.0 nC. Find the electric field (magnitude and direction) at a distance (a) r 1.5 m (b) r2 4.0 m from the center. (c) Find the charges on the inner and outer surfaces of the spherical shell
Answer #1

from Gauss law,

E (4 pi r^2) = Qin/e0


(A) r1 = 1.5 m

Qin = q1

E ( 4 x pi x 1.5^2) = (9 x 10^-9) / (8.854 x 10^-12)

E = 35.95 N/C radially outward  

(B) r2 is inside the metal so E = 0

{ field inside conductor is always zero.}


(C) Qinner + Q1 = 0

Qinner = - q1 = - 9 nC


Qinner + Qouter = Q2 = - 6 nC

- 9nC + Qouter = - 6

Qouter = 3 nC